Assume you buy a bond with the following features Bond maturity = 6 Coupon Rate = 4.00% Face Value = $1,000 Annual Coupons When you buy the bond the market interest rate = 4.00% Immediately after you buy the bond the interest rate changes to 3.50% What is the "price risk" effect in year 3 ? a) $14.43 b) -$14.01 c) -$14.43 d) $14.01 e) -$13.59 f) $13.59
Answer #1

Interest rate 4%:

Price = 4%

When coupon rate is equal to interest rate, bonds will always sell at apr.

Interest rate 3.50%

Coupon = 0.04 * 1000 = 40

Price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 +r)n

Price = 40 * [1 - 1 / (1 + 0.035)3] / 0.035 + 1000 / (1 +0.035)3

Price = 40 * [1 - 0.901943] / 0.035 + 901.942706

Price = 40 * 2.801637 + 901.942706

Price = $1,014.01

Change in price = $1,014.01 - 1,000

Change in price = $14.01

Price risk affect will be an increase in price by $14.01